Solve using matrices
2x-3y=4
5x+4y = 33

Answer by Hook(36)

if we put this into a matrix, we get:
[2 -3 4]
[5 4 33]
I'm going to multiply the top line by 5 and the bottom line by -2
[10 -15 20]
[-10 -8 -66]
I'll at the top and bottom lines together and make the sum be the new bottom line
[10 -15 20]
[ 0 -23 -46]
ok...I'm going to get back to some smaller numbers. I'll divide the top line by 5 and the bottom line by -23
[2 -3 4]
[0 1 2]
Now, I'll multipy the bottom line by 3
[2 -3 4]
[0 3 6]
Here, I'll add the two lines together and put them in the top line
[2 0 10]
[0 3 6]
Almost there. I'm going to get to the identity matrix by dividing the top line by 2 and the bottom line by 3
[1 0 5]
[0 1 2]
We get x = 5 and y = 2!
CHECK:
We'll plug our answers into the original equation to check our work
2x-3y=4
2(5) - 3(2) = 4
10 - 6 = 4
4 =4 Check!
5x+4y = 33
5(5) + 4(2) = 33
25 + 5 = 33
33 = 33 Check!